# On the multiplier spectrum of polynomials

### v1

@@ 3.2. The case of degree $2\leq d\leq 3$
By Lemma 3.27 and Theorems 3.25 and 3.26, there exists a non-empty Zariski open subset $U$ of ${\rm Poly}^{d}(\mathbb{C})$ such that every $F\in U$ is not a generalized Lattès map and satisfies conditions (1)–(3) of Theorem 3.25 (for $d=2$) or Theorem 3.26 (for $d=3$).
Let $F\in U$, and suppose $B,X\in\mathbb{C}(z)\setminus\mathbb{C}$ satisfy $F^{\circ k}\circ X=X\circ B$ for some $k\in\mathbb{Z}_{>0}$. By Remark 3.19, $F^{\circ k}$ is not a generalized Lattès map. Applying [Pak23, Proposition 3.3], there exist $Y\in\mathbb{C}(z)\setminus\mathbb{C}$ and $r\in\mathbb{Z}_{\geq 0}$ such that $X\circ Y=F^{\circ kr}$. By (3) of Theorem 3.25 or 3.26, $X=F^{\circ m}\circ\delta$ for some $m\in\mathbb{Z}_{\geq 0}$ and $\delta\in{\rm PGL}_{2}(\mathbb{C})$. Then
(3.6) $F^{\circ(k+m)}=F^{\circ m}\circ\delta\circ B\circ\delta^{-1}.$
Again by (3), there exists $\nu\in{\rm PGL}_{2}(\mathbb{C})$ such that $\delta\circ B\circ\delta^{-1}=\nu^{-1}\circ F^{\circ k}$. Then $F^{\circ(k+m)}=F^{\circ m}\circ\nu^{-1}\circ F^{\circ k}$, so $F^{\circ m}=F^{\circ m}\circ\nu$ and $\nu\in\Sigma_{\infty}(F)$ (if $m=0$, then $\nu(z)=z\in\Sigma_{\infty}(F)$). By (1), $\nu\in\Sigma(F)$. Let $\mu:=\nu\circ\delta$. Then
$X=F^{\circ m}\circ\delta=F^{\circ m}\circ\nu\circ\delta=F^{\circ m}\circ\mu,$
@@
$B=\delta^{-1}\circ\nu^{-1}\circ F^{\circ k}\circ\delta=\delta^{-1}\circ\nu^{-1}\circ F^{\circ k}\circ\nu\circ\delta=\mu^{-1}\circ F^{\circ k}\circ\mu.$
@@
Let $d\in\{2,3\}$. There exists a non-empty Zariski open subset $U$ of ${\rm Poly}^{d}(\mathbb{C})$ such that for all $(F_{1},F_{2})\in U\times U$ and every irreducible algebraic curve $C\subset\mathbb{P}^{1}_{\mathbb{C}}\times\mathbb{P}^{1}_{\mathbb{C}}$ that is not a horizontal or vertical line, the following are equivalent:
  1. (1)
@@
  2. (2)
    there exist $\alpha\in{\rm PGL}_{2}(\mathbb{C})$ and $s\in\mathbb{Z}_{\geq 0}$ such that $F_{2}=\alpha\circ F_{1}\circ\alpha^{-1}$ and $C$ is one of the graphs y=(αF_1^s)(x) or x=(F_1^sα^-1)(y), where $x$ and $y$ are the coordinates on $\mathbb{P}^{1}_{\mathbb{C}}\times\mathbb{P}^{1}_{\mathbb{C}}$.
@@
By Lemma 3.27 and Theorems 3.25 and 3.26, there exists a non-empty Zariski open subset $U$ of ${\rm Poly}^{d}(\mathbb{C})$ such that every $F\in U$ is not a generalized Lattès map and satisfies (1)–(3) of Theorem 3.25 (for $d=2$) or 3.26 (for $d=3$). By the proof of Theorem 3.28, the conclusion of that theorem holds for every $F\in U$.
Let $F_{1},F_{2}\in U$ have the same degree $d$, and let $C\subset\mathbb{P}^{1}_{\mathbb{C}}\times\mathbb{P}^{1}_{\mathbb{C}}$ be an irreducible curve that is not a vertical or horizontal line.
Assume (1): $C$ is $(F_{1},F_{2})$-periodic. Then for some $k\in\mathbb{Z}_{>0}$, $C$ is $(F_{1}^{\circ k},F_{2}^{\circ k})$-invariant. By Remark 3.19, both $F_{1}^{\circ k}$ and $F_{2}^{\circ k}$ are not a generalized Lattès map. Applying [Pak23, Theorem 1.1], there exist non-constant rational maps $X_{1},X_{2}\in\mathbb{C}(z)$ such that:
@@
By Theorem 3.28, there exist $\alpha_{1},\alpha_{2}\in{\rm PGL}_{2}(\mathbb{C})$ and $m_{1},m_{2}\in\mathbb{Z}_{\geq 0}$ such that
@@
so $F_{2}=\alpha\circ F_{1}\circ\alpha^{-1}$ by (2) of Theorem 3.25 or 3.26. The curve $C$ is parametrized by
$t\mapsto(X_{1}(t),X_{2}(t))=(F_{1}^{\circ m_{1}}\circ\alpha_{1}(t),\alpha\circ F_{1}^{\circ m_{2}}\circ\alpha_{1}(t)),$
so $t\mapsto(F_{1}^{\circ m_{1}}(t),\alpha\circ F_{1}^{\circ m_{2}}(t))$ also parametrizes $C$.
If $m_{1}\geq m_{2}$, set $s=m_{1}-m_{2}\in\mathbb{Z}_{\geq 0}$; then $t\mapsto(F_{1}^{\circ s}\circ\alpha^{-1}(t),t)$ parametrizes $C$. If $m_{1}<m_{2}$, set $s=m_{2}-m_{1}\in\mathbb{Z}_{>0}$; then $t\mapsto(t,\alpha\circ F_{1}^{\circ s}(t))$ parametrizes $C$. Thus, $C$ is one of the stated graphs.
Conversely, if (2) holds, then $C$ is clearly $(F_{1},F_{2})$-invariant. ∎
@@
Compared to Theorem 3.22, the form of periodic curves in Theorem 3.29 is simpler due to Theorems 3.25(2) and 3.26(2). By (2) of Theorem 3.25 or 3.26, for two general polynomials $F_{1},F_{2}\in\mathbb{C}[z]$ of the same degree $d\in\{2,3\}$, $F_{1}$ and $F_{2}$ are conjugate if and only if $F_{1}^{\circ k}$ and $F_{2}^{\circ k}$ are conjugate for some $k\in\mathbb{Z}_{>0}$. If $F$ satisfies (2) of Theorem 3.25 or 3.26, then ${\rm Aut}(F)={\rm Aut}_{\infty}(F)$. For $\nu\in{\rm Aut}(F)$, we have $\nu\circ F\circ\nu^{-1}=F$, so we may replace $\alpha$ in Theorem 3.22 by $\alpha\circ\nu$ or $\alpha\circ\nu^{-1}$ when $\nu\in{\rm Aut}(F_{1})$.
Proposition 3.31 (Proposition 1.6).
Let $d\in\{2,3\}$. There exists a non-empty Zariski open subset $U$ of ${\rm Poly}^{d}(\mathbb{C})$ such that for all $(f,g)\in U\times U$ with $f$ and $g$ intertwined, we have $[f]=[g]$ in ${\rm MPoly}^{d}(\mathbb{C})$.
@@
Using Theorem 3.23 or Proposition 3.31 in place of [FG22, Theorems 3.51 and 3.52], we obtain a proof of Theorem 1.4 following Ji-Xie [JX25].
@@
Let $d\geq 2$ be an integer. Assume by contradiction that $\tilde{\tau}_{d}$ is not generically injective. By Theorem 3.23 or Proposition 3.31, there exists a non-empty Zariski open subset $U$ of ${\rm MPoly}^{d}(\mathbb{C})$ such that for all $f,g\in{\rm Poly}^{d}(\mathbb{C})$ with $f],[g]\in U$,
@@
Take a non-empty Zariski open subset $W$ of the Zariski closure of $\tilde{\tau}_{d}(U)$ such that $\tilde{\tau}_{d}^{-1}(W)\subseteq U$ and $\tilde{\tau}_{d}:\tilde{\tau}_{d}^{-1}(W)\to W$ is a finite étale morphism of degree at least $2$. After shrinking $U$, we may assume $U=\tilde{\tau}_{d}^{-1}(W)$.
It is well known that the PCF locus $\{f\in{\rm Poly}^{d}(\mathbb{C}):f\text{ is }{\rm PCF}\}$ is Zariski-dense in ${\rm Poly}^{d}$. As in [JX25, Proof of Theorem 1.3], we can find two non-isotrivial algebraic families $h_{1},h_{2}$ of degree $d$ polynomials parameterized by an irreducible affine curve $C$ with $h_{1}(C)\cup h_{2}(C)\subseteq U$ and
$\tilde{\tau}_{d}\circ h_{1}=\tilde{\tau}_{d}\circ h_{2}:C\to\mathbb{A}_{\mathbb{C}}^{N}$
such that for every $t\in C$, $h_{1}(t)\neq h_{2}(t)$ in ${\rm MPoly}^{d}$, and the PCF locus
$\{t\in C:h_{1}(t)\text{ is }{\rm PCF}\}$
for $h_{1}$ is Zariski-dense in $C$. Here $N=N_{d,1}+\cdots N_{d,m_{d}}$.
By [JXZ25, Theorem 1.14], whether a rational map is PCF is determined by its length spectrum (hence its multiplier spectrum). So
$\{t\in C:h_{2}(t)\text{ is }{\rm PCF}\}=\{t\in C:h_{1}(t)\text{ is }{\rm PCF}\},$
which is Zariski-dense in $C$. Applying the variant of DAO (Dynamical André-Oort) conjecture for curves proved in [JX25, Theorem 3.4], the set
$\{t\in C:h_{1}(t)\text{ and }h_{2}(t)\text{ are intertwined}\}$
is infinite. (The condition that $h_{1},h_{2}$ are not families of flexible Lattès maps holds automatically for polynomials.) Then by (3.7), the set
$\{t\in C:h_{1}(t)=h_{2}(t)\text{ in }{\rm MPoly}^{d}\}$
is infinite, contradicting the construction of $h_{1}$ and $h_{2}$. ∎
3.4. A case of the Zariski-dense orbit conjecture
As a byproduct, we now present a proof of the Zariski-dense orbit conjecture (ZDO) for a general split polynomial endomorphism on $(\mathbb{P}^{1})^{2}$ with all factors of the same degree $d\geq 2$, which is essentially due to Pakovich [Pak25a, Pak25b].
Conjecture 3.32 (ZDO).
Let $k$ be an algebraically closed field of characteristic 0. Given an irreducible quasi-projective variety $X$ over $k$ and a dominant rational self-map $f$ on $X$. If $\{g\in k(X):g\circ f=g\}=k$, where $k(X)$ is the function field of $X$, then there exists $x\in X(k)$ whose forward orbit under $f$ is well-defined and Zariski-dense in $X$.
@@ 3.4. A case of the Zariski-dense orbit conjecture
Fix an integer $d\geq 2$. For a general pair of polynomials $F_{1},F_{2}\in{\rm Poly}^{d}(\mathbb{C})$, the endomorphism $(F_{1},F_{2}):\mathbb{P}^{1}_{\mathbb{C}}\times\mathbb{P}^{1}_{\mathbb{C}}\to\mathbb{P}^{1}_{\mathbb{C}}\times\mathbb{P}^{1}_{\mathbb{C}}$ has no irreducible periodic curves other than vertical or horizontal lines. In particular, for every $x\in\mathbb{P}^{1}(\mathbb{C})\setminus{\rm PrePer}(F_{1})$ and $y\in\mathbb{P}^{1}(\mathbb{C})\setminus{\rm PrePer}(F_{2})$, the Zariski closure of the forward orbit $\mathcal{O}_{(F_{1},F_{2})}((x,y))$ is $\mathbb{P}^{1}_{\mathbb{C}}\times\mathbb{P}^{1}_{\mathbb{C}}$.
@@
For $d\in\{2,3\}$, the result follows from Theorem 3.29.
Assume $d\geq 4$. By Lemmas 3.15 and 3.16, a general $f\in{\rm Poly}^{d}(\mathbb{C})$ is pre-simple with $G(f)=1$. Let $F_{1},F_{2}$ be pre-simple polynomials of degree $m$ with $G(F_{1})=1$, $G(F_{2})=1$, and $F_{1}]\neq[F_{2}]$. Such pairs are general in ${\rm Poly}^{d}(\mathbb{C})\times{\rm Poly}^{d}(\mathbb{C})$.
Suppose $C$ is an irreducible $(F_{1},F_{2})$-periodic curve in $\mathbb{P}^{1}_{\mathbb{C}}\times\mathbb{P}^{1}_{\mathbb{C}}$ that is not a vertical or horizontal line. By Theorem 3.22, $F_{2}^{\circ k}=\alpha\circ F_{1}^{\circ k}\circ\alpha^{-1}$ for some $k\in\mathbb{Z}_{>0}$ and $\alpha\in{\rm PGL}_{2}(\mathbb{C})$. By Theorem 3.14 and $G_{0}(F_{2})\subseteq G(F_{2})=1$, we have $E(F_{2})=\left\langle F_{2}\right\rangle$. Since
@@
we have $\alpha\circ F_{1}\circ\alpha^{-1}\in E(F_{2})=\left\langle F_{2}\right\rangle$. Comparing degrees, $\alpha\circ F_{1}\circ\alpha^{-1}=F_{2}$, contradicting $F_{1}]\neq[F_{2}]$. ∎
@@
Xie [Xie17] proved ZDO for all dominant polynomial endomorphisms $f:\mathbb{A}_{k}^{2}\to\mathbb{A}_{k}^{2}$ over an algebraically closed field $k$ of characteristic 0 using valuative techniques.
4. Polynomials with the same multiplier spectrum

### v2

@@ 3.2. The case of degree $2\leq d\leq 3$
By Lemma 3.33 and Theorems 3.31 and 3.32, there exists a non-empty Zariski open subset $U$ of $\mathrm{Poly}^{d}(\mathbb{C})$ such that every $F\in U$ is not a generalized Lattès map and satisfies $\Sigma(F)=\Sigma_{\infty}(F)$.
Let $F\in U$, and let $B,X\in\mathbb{C}(z)\setminus\mathbb{C}$ be such that $F^{\circ k}\circ X=X\circ B$ for some $k\in\mathbb{Z}_{>0}$. If $\deg(X)=1$, then the conclusion trivially holds with $(m,\mu)=(0,X)$. We may assume $\deg(X)\geq 2$. By Remark 3.25, $F^{\circ k}$ is not a generalized Lattès map. Applying [Pak23, Proposition 3.3], there exist $Y\in\mathbb{C}(z)\setminus\mathbb{C}$ and $r\in\mathbb{Z}_{>0}$ such that $X\circ Y=F^{\circ kr}$. By Corollary 3.12, $m:=\log_{d}\deg(X)$ is a positive integer and $X=F^{\circ m}\circ\delta$ for some $\delta\in\mathrm{PGL}_{2}(\mathbb{C})$. Then
(3.9) $F^{\circ(k+m)}=F^{\circ k}\circ F^{\circ m}=F^{\circ k}\circ X\circ\delta^{-1}=X\circ B\circ\delta^{-1}=F^{\circ m}\circ\delta\circ B\circ\delta^{-1}.$
Again by Corollary 3.12, there exists $\nu\in\mathrm{PGL}_{2}(\mathbb{C})$ such that $\delta\circ B\circ\delta^{-1}=\nu^{-1}\circ F^{\circ k}$, hence $F^{\circ(k+m)}=F^{\circ m}\circ\nu^{-1}\circ F^{\circ k}$ and $F^{\circ m}=F^{\circ m}\circ\nu$. Then $\nu\in\Sigma_{\infty}(F)=\Sigma(F)$. Let $\mu:=\nu\circ\delta$. Then
$X=F^{\circ m}\circ\delta=(F^{\circ m}\circ\nu)\circ\delta=F^{\circ m}\circ\mu,$
@@
$B=\delta^{-1}\circ\nu^{-1}\circ F^{\circ k}\circ\delta=\mu^{-1}\circ(F^{\circ k}\circ\nu)\circ\delta=\mu^{-1}\circ F^{\circ k}\circ\mu.$
@@
Let $d\in\{2,3\}$. There exists a non-empty Zariski open subset $U$ of $\mathrm{Poly}^{d}(\mathbb{C})$ such that for all $(F_{1},F_{2})\in U\times U$ and every irreducible algebraic curve $C\subset\mathbb{P}^{1}\times\mathbb{P}^{1}$ over $\mathbb{C}$ that is neither a horizontal nor a vertical line, the following are equivalent:
  1. (i)
@@
  2. (ii)
    there exist $s\in\mathbb{Z}_{\geq 0}$ and a degree one polynomial $\alpha\in\mathrm{Aff}(\mathbb{C})$ such that F_2=αF_1α^-1 and $C$ is one of the graphs y=(αF_1^s)(x) or x=(F_1^sα^-1)(y), where $x$ and $y$ are the coordinates on $\mathbb{P}^{1}\times\mathbb{P}^{1}$.
@@
By Lemma 3.33 and Theorems 3.313.32, and 3.34, there exists a non-empty Zariski open subset $U$ of $\mathrm{Poly}^{d}(\mathbb{C})$ such that every $F\in U$ is not a generalized Lattès map, satisfies properties (1) and (2) of Theorem 3.31 (for $d=2$) or Theorem 3.32 (for $d=3$), and satisfies the conclusion of Theorem 3.34.
Let $(F_{1},F_{2})\in U^{2}$, and let $C\subset\mathbb{P}^{1}\times\mathbb{P}^{1}$ be an irreducible curve over $\mathbb{C}$ that is neither a horizontal nor a vertical line.
Assume (i): $C$ is $(F_{1},F_{2})$-periodic. Fix $k\in\mathbb{Z}_{>0}$ such that $C$ is $(F_{1}^{\circ k},F_{2}^{\circ k})$-invariant. By Remark 3.25, neither $F_{1}^{\circ k}$ nor $F_{2}^{\circ k}$ is a generalized Lattès map. By [Pak23, Theorem 1.1], there exist $X_{1},X_{2},B\in\mathbb{C}(z)\setminus\mathbb{C}$ such that:
@@
By Theorem 3.34, there exist $\alpha_{1},\alpha_{2}\in\mathrm{PGL}_{2}(\mathbb{C})$ and $m_{1},m_{2}\in\mathbb{Z}_{\geq 0}$ such that
@@
so $F_{2}=\alpha\circ F_{1}\circ\alpha^{-1}$ by (2) of Theorem 3.31 or 3.32. For $1\leq i\leq 2$, since $F_{i}$ is not a generalized Lattès map (in particular, not conjugate to $z^{d}$), $\infty$ is the unique totally invariant fixed point of $F_{i}$. Hence, the Möbius transformation $\alpha$ must fix $\infty$, so $\alpha\in\mathrm{Aff}(\mathbb{C})$.
The curve $C$ is also parametrized by $t\mapsto(F_{1}^{\circ m_{1}}(t),\alpha\circ F_{1}^{\circ m_{2}}(t))$. If $m_{1}\geq m_{2}$, then $t\mapsto(F_{1}^{\circ s}\circ\alpha^{-1}(t),t)$ parametrizes $C$, where $s:=m_{1}-m_{2}\in\mathbb{Z}_{\geq 0}$. If $m_{1}<m_{2}$, then $t\mapsto(t,\alpha\circ F_{1}^{\circ s}(t))$ parametrizes $C$, where $s:=m_{2}-m_{1}\in\mathbb{Z}_{>0}$. Thus, $C$ is one of the graphs in (ii).
Conversely, if (ii) holds, then $C$ is clearly $(F_{1},F_{2})$-invariant. ∎
@@
Let $d\in\{2,3\}$. Compared to Theorem 3.28, the form of periodic curves in Theorem 3.35 is relatively simple. This is mainly due to (2) of Theorem 3.31 or 3.32, which implies:
  • for two general polynomials $F_{1},F_{2}\in\mathrm{Poly}^{d}(\mathbb{C})$, $F_{1}$ and $F_{2}$ are conjugate if and only if $F_{1}^{\circ k}$ and $F_{2}^{\circ k}$ are conjugate for some $k\in\mathbb{Z}_{>0}$;
  • for a general polynomial $F\in\mathrm{Poly}^{d}(\mathbb{C})$, $\mathrm{Aut}(F)=\mathrm{Aut}_{\infty}(F)$.
Proposition 3.37 (=Proposition 1.7).
Let $d\in\{2,3\}$. There exists a non-empty Zariski open subset $U$ of $\mathrm{Poly}^{d}(\mathbb{C})$ such that for all $(f,g)\in U\times U$, $f$ and $g$ are intertwined if and only if $[f]=[g]$ in $\mathrm{MPoly}^{d}(\mathbb{C})$.
@@
Using Theorem 3.29 or Proposition 3.37 in place of [FG22, Theorems 3.51 and 3.52], we obtain a proof of Theorem 1.4 following Ji and Xie [JX25].
@@
Let $d\geq 2$ be an integer. Suppose, for the sake of contradiction, that $\tilde{\tau}_{d}$ is not generically injective. By Theorem 3.29 or Proposition 3.37, there exists a non-empty Zariski open subset $U$ of $\mathrm{MPoly}^{d}(\mathbb{C})$ such that for all $f,g\in\mathrm{Poly}^{d}(\mathbb{C})$ with $f],[g]\in U$,
@@
Let $W$ be a non-empty Zariski open subset of the Zariski closure of $\tilde{\tau}_{d}(U)$ such that $\tilde{\tau}_{d}^{-1}(W)\subseteq U$ and $\tilde{\tau}_{d}:\tilde{\tau}_{d}^{-1}(W)\to W$ is a finite étale morphism of degree at least $2$. After shrinking $U$, we may assume $U=\tilde{\tau}_{d}^{-1}(W)$.
It is well-known that the PCF locus $\{f\in\mathrm{Poly}^{d}(\mathbb{C})\colon f\text{ is PCF}\}$ is Zariski-dense in $\mathrm{Poly}^{d}(\mathbb{C})$. As in the proof of [JX25, Theorem 1.3], we can find two non-isotrivial algebraic families $h_{1},h_{2}$ of degree $d$ polynomials parametrized by an irreducible affine curve $C$ with $h_{i}(C(\mathbb{C}))\subseteq U$ for $1\leq i\leq 2$ and
$\tilde{\tau}_{d}\circ h_{1}=\tilde{\tau}_{d}\circ h_{2}\colon C\to\mathbb{A}_{\mathbb{C}}^{N}$
(where $N=N_{d,1}+\cdots+N_{d,m_{d}}\in\mathbb{Z}_{>0}$) such that
(3.11) $\text{for every }t\in C(\mathbb{C})\text{, }h_{1}(t)\neq h_{2}(t)\text{ in }\mathrm{MPoly}^{d}\text{,}$
and the PCF locus
$\{t\in C(\mathbb{C})\colon h_{1}(t)\text{ is PCF}\}$
of $h_{1}$ is Zariski-dense in $C$.
By [JXZ25, Theorem 1.14], the property of being PCF is determined by the length spectrum (in particular, by the multiplier spectrum). So
$\{t\in C(\mathbb{C})\colon h_{2}(t)\text{ is PCF}\}=\{t\in C(\mathbb{C})\colon h_{1}(t)\text{ is PCF}\},$
which is Zariski-dense in $C$. By the DAO-type theorem [JX25, Theorem 1.5] proved by Ji and Xie, we deduce that
$\{t\in C(\mathbb{C})\colon h_{1}(t)\text{ and }h_{2}(t)\text{ are intertwined}\}=\infty.$
Note that polynomials of degree $\geq 2$ cannot be Lattès, since the Julia set of any Lattès map is the entire Riemann sphere [Mil06b, § 2]. By (3.10), we obtain
$\{t\in C(\mathbb{C})\colon h_{1}(t)=h_{2}(t)\text{ in }\mathrm{MPoly}^{d}\}=\infty,$
contradicting (3.11). ∎
3.4. General polynomial pairs and the Zariski-dense orbit conjecture
As a byproduct, we present a proof of the Zariski-dense orbit conjecture for a general split polynomial endomorphism on $\mathbb{A}^{2}$ with all factors of the same degree $d\geq 2$, which is essentially due to Pakovich [Pak25a, Pak25b].
Conjecture 3.38 (Zariski-dense orbit conjecture = ZDO).
Let $k$ be an algebraically closed field of characteristic zero. Let $X$ be an irreducible quasi-projective variety over $k$ and $f$ a dominant rational self-map on $X$. If $\{g\in k(X)\colon g\circ f=g\}=k$, where $k(X)$ is the function field of $X$, then there exists $x\in X(k)$ whose forward orbit under $f$ is well-defined and Zariski-dense in $X$.
We remark that the converse of ZDO holds and is easy to prove.
@@ 3.4. General polynomial pairs and the Zariski-dense orbit conjecture
Fix an integer $d\geq 2$. For a general pair of polynomials $(F_{1},F_{2})\in\mathrm{Poly}^{d}(\mathbb{C})^{2}$, the endomorphism $(F_{1},F_{2}):\mathbb{A}^{2}\to\mathbb{A}^{2}$ has no irreducible periodic curves other than vertical or horizontal lines. In particular, for all $x\in\mathbb{A}^{1}(\mathbb{C})\setminus\mathrm{PrePer}(F_{1})$ and $y\in\mathbb{A}^{1}(\mathbb{C})\setminus\mathrm{PrePer}(F_{2})$, the forward orbit $O_{(F_{1},F_{2})}((x,y))$ is Zariski-dense in $\mathbb{A}^{2}$.
@@
Observe that for a general pair of polynomials $(F_{1},F_{2})\in\mathrm{Poly}^{d}(\mathbb{C})^{2}$, we have $F_{1}]\neq[F_{2}]$ in $\mathrm{MPoly}^{d}$. For $d\in\{2,3\}$, the result follows from Theorem 3.35.
Assume $d\geq 4$. Let $F_{1},F_{2}$ be pre-simple polynomials of degree $d$ such that
$G(F_{1})=G(F_{2})=1\quad\text{and}\quad[F_{1}]\neq[F_{2}].$
By Lemmas 3.21 and 3.22, such pairs are general in $\mathrm{Poly}^{d}(\mathbb{C})\times\mathrm{Poly}^{d}(\mathbb{C})$.
Suppose that $C\subset\mathbb{A}^{2}$ is an irreducible $(F_{1},F_{2})$-periodic curve over $\mathbb{C}$ that is neither a horizontal nor a vertical line. By Theorem 3.28, $F_{2}^{\circ k}=\alpha\circ F_{1}^{\circ k}\circ\alpha^{-1}$ for some $k\in\mathbb{Z}_{>0}$ and $\alpha\in\mathrm{Aff}(\mathbb{C})$. By Theorem 3.20 and the fact that $G_{0}(F_{2})\subseteq G(F_{2})=1$, we have $E(F_{2})=\left\langle F_{2}\right\rangle$. Since
@@
we see that $\alpha\circ F_{1}\circ\alpha^{-1}\in E(F_{2})=\left\langle F_{2}\right\rangle$. Comparing degrees gives $\alpha\circ F_{1}\circ\alpha^{-1}=F_{2}$, contradicting $F_{1}]\neq[F_{2}]$. ∎
@@
Let $F_{i}\in\mathbb{C}[z]$ be a polynomial of degree $d_{i}\geq 2$ for $1\leq i\leq 2$ such that $d_{1}\neq d_{2}$. In the non-equal-degree case, $F_{1}$ and $F_{2}$ are not intertwined (see Remark 1.6), so the endomorphism $(F_{1},F_{2}):\mathbb{A}^{2}\to\mathbb{A}^{2}$ has no irreducible periodic curves other than vertical or horizontal lines.
Remark 3.41.
Note that for any polynomial $F(z)\in k[z]$ of degree $d\geq 2$ over a field $k$ of characteristic zero, we have $\mathbb{A}^{1}(k)\setminus\mathrm{PrePer}(F_{1})\neq\emptyset$. It is easy to see that for a general pair $(F_{1},F_{2})$ of polynomials of degree $d\geq 2$ defined over $\overline{\mathbb{Q}}$, the analogous conclusion holds over $\overline{\mathbb{Q}}$.
Remark 3.42.
Xie [Xie17] proved ZDO for all dominant polynomial endomorphisms $f:\mathbb{A}_{k}^{2}\to\mathbb{A}_{k}^{2}$ over an algebraically closed field $k$ of characteristic zero using valuative techniques.
4. Isospectral polynomials